# Generalised Fermat equation: a survey of solved cases

### v1

@@ 4.1 The case $a=b=1$
We start with the case $a=b=1$. In this instance equation (4.1) reduces to
@@
We remark that for equation (4.2) Proposition 1.3 is applicable if and only if $c$ is not divisible by a $p$-th power of any prime. If this condition is satisfied, we may talk about solutions in pairwise coprime integers $(x,y,z)$. From symmetry, we may also assume that $x>y$.
In 2023, Freitas and Najman [71] considered the equation (4.2) where $p,r>3$ are primes and $c$ is an odd positive integer, not an $r$-th power and $r\nmid c$. They obtained that for a set of prime exponents $p$ of positive density, the equation has no non-trivial primitive solutions $(x,y,z)$ such that $2\mid x+y$ or $r\mid x+y$.
In 2024, Kara, Mocanu and Özman [85] prove some results conditional on the weak Frey-Mazur conjecture [85, Conjecture 3.9] and the Eichler-Shimura conjecture [74, 85, Conjecture 3.11].
@@
    If $p\equiv 3$ modulo $4$, then there exists a (non-effective) constant ${\cal{B}}_{p}$ such that for any prime $r$ with $r>{\cal{B}}_{p}$, the equation (4.2) has no non-trivial primitive integer solutions.
@@
    If $p\equiv 1$ modulo $4$, then there exists a (non-effective) constant ${\cal{B}}^{\prime}_{p}$ such that for any prime $r$ with $r>{\cal{B}}^{\prime}_{p}$, the equation (4.2) has no non-trivial primitive integer solutions $(x,y,z)$ where $p|z$.
@@
Let $p>5$ be a fixed prime such that $p\equiv 1$ modulo $4$ and $c$ be a non-zero integer with $\gcd(c,p)=1$. Assume the Weak Frey-Mazur Conjecture and the Eichler-Shimura Conjecture. Then there exists a (non-effective) constant ${\cal{B}}^{\prime\prime}_{p}$ such that for any prime $r$ with $r>{\cal{B}}^{\prime\prime}_{p}$, the equation (4.2) has no non-trivial primitive integer solutions.
4.1.1 Equation (4.2) with fixed $p$
In this section, we consider equation (4.2) with $p\geq 3$ fixed and $r$ varies. In this section, we will not assume that $p$ is a prime. We start with the case $p=3$.
The equation (4.2) with $p=3$, $c=2$ and $r$ even was solved by Zhang and Bai [148].
@@
The equation (4.2) with $p=3$, $c=s^{\alpha}$ with $s$ prime, $\alpha\geq 1$ and $r\geq 4$, that is,
@@
There is an explicit set $T$ of primes containing a positive proportion of all primes [^15] Specifically, $T$ is the set of primes $s$ for which there does not exist an elliptic curve with rational 2-torsion and conductor $2^{M}3^{2}s$, $1\leq M\leq 3$., such that for every $s\in T$, every integer $\alpha\geq 1$, and every prime $r$ satisfying $r\geq s^{8s}$ and $r\nmid\alpha$, the equation (4.3) has no solutions in coprime non-zero integers $x,y,z$.
@@
There exists a set ${\cal W}_{3}$ of primes, with #$\#\{s\leq x:s\in{\cal W}_{3}\}\leq K\sqrt{x}\log^{2}(x)$ for some $K>0$, such that for every prime $s\geq 5$ with $s\not\in{\cal W}_{3}$, any $\alpha\geq 1$, and every prime $r$ satisfying $r\geq s^{2s}$, equation (4.3) has no solutions in coprime non-zero integers $x,y,z$.
Because the number of all primes up to $x$ is about $\frac{x}{\log x}$, set ${\cal W}_{3}$ contains a negligible part of all primes, hence Theorem 4.5 solves equation (4.3) for almost all primes $s$. There is an algorithm for checking whether any particular prime belongs to ${\cal W}_{3}$. For example, it is known [11] that all primes $s$ equal to $317$ or $1757$ modulo $2040$ do not belong to ${\cal W}_{3}$, hence Theorem 4.5 is applicable to all such primes.
Theorem 4.5 is not applicable for primes $s\in{\cal W}_{3}$. In 2018, Bennett, Bruni and Freitas [5] proved a more general version of this theorem with a smaller exceptional set.
@@
For every prime $s\geq 5$ such that $s\not\in S_{0}$, any $\alpha\geq 1$, and any prime $r$ satisfying $r\geq s^{2s}$, equation (4.3) has no solutions in coprime non-zero integers $x,y,z$.
As mentioned in [5], $S_{0}\subset{\cal W}_{3}$, hence Theorem 4.6 is a generalization of Theorem 4.5. The smallest examples of primes in ${\cal W}_{3}\setminus S_{0}$ are
$s=53,83,149,167,173,199,223,227,233,263,281,293,311,347,353,359,389,\dots$
@@
where $s$ and $r$ are primes. For small primes $s\in S_{0}$, they determined [5, Theorem 7.2] the modularity conditions on $r$ such that for $r\geq s^{2s}$ the equation has no primitive solutions, see Table 4.1.
@@
$5$ $13,19$ or $23$ (mod 24) $47$ $5,11,13,17,19$ or $23$ (mod 24)
$11$ $13,17,19$ or $23$ (mod 24) $95$,7,11,13,19571113195,7,11,13,195 , 7 , 11 , 13 , 19 or $23$ (mod 24)
$13$ $11$ (mod 12) $67$ $7,11,13,29,37,41,43,59,67,71,89,101$ or $103$ (mod 120)
$17$ $5,17$ or $23$ (mod 24) $71$ $5$ (mod 6)
$23$ $19$ or $23$ (mod 24) $73$ $41,71$ or $89$ (mod 120)
$29$ $7,11,13,17,19$ or $23$ (mod 24) $79$ $5,7,11,13,19$ or $23$ (mod 24)
$31$ $5$ or $11$ (mod 24) $89$ $13,17,19$ or $23$ (mod 24)
$41$ $5,7,11,17,19$ or $23$ (mod 24) $97$ $11$ (mod 12)
Table 4.1: Modularity conditions on $r$ for small primes $s$, such that the equation (4.5) has no primitive integer solutions if $r\geq s^{2s}$.
@@
If $r$ is prime with $r\equiv 13,19$ or $23$ (mod $24$), then there are no non-trivial primitive solutions $(x,y,z)$ to the equation
@@
Set $S_{0}$ contains all primes of the form $s=2^{a}3^{b}-1$ for integers $a\geq 5$ and $b\geq 1$, hence these primes are not covered by Theorem 4.6. For such primes, we have the following result [5], which uses a Jacobi symbol condition.
@@
then the equation (4.3) has no non-trivial primitive solutions.
@@
We next discuss equation (4.2) with $p=4$ and $c,r$ primes, that is
@@
We first discuss equation (4.7) with $c=73,89$ or $113$ and $r>13$. Dieulefait [63] proved the following theorem.
@@
If $r$ is a prime with $r>13$, then the equation (4.7) with $c=73,89$ or $113$ does not have any non-trivial primitive integer solutions.
@@
We next discuss the equation (4.2) with $p=5$, that is,
@@
There exists a set ${\cal W}_{5}$ of primes, with #$\#\{s\leq x:s\in{\cal W}_{5}\}\leq K\sqrt{x}\log^{2}(x)$ for some $K>0$, such that for every prime $s\geq 5$ such that $s\not\in{\cal W}_{5}$, any $\alpha\geq 1$, and any prime $r$ satisfying $r\geq s^{13s}$, equation
@@
Bruni [33] also developed methods for solving equation (4.9) for certain primes $s$ belonging to ${\cal W}_{5}$. In fact, he solved equation (4.9) for all primes $s$ outside the specific families listed in Table 7.6 of [33].
For the equation (4.8) with $c=2$, that is,
@@
For all primes $r$, the equation (4.10) has no non-trivial primitive solutions $(x,y,z)$ such that $z$ is divisible by either $2$, $5$, or $r$.
Equation (4.8) with $c=3$, that is,
@@
Equation (4.11) has no non-trivial primitive integer solutions for any integer $r\geq 2$.
We next discuss equation (4.8) for other values of $c$. In 2013, Dieulefait and Freitas [59] proved the following theorem.
@@
    For any prime $r>13$ such that $r\equiv 1$ (mod 4), or $r\equiv\pm 1$ (mod 5), the equation (4.8) with $c=2s$ has no non-trivial solutions with $\gcd(x,y)=1$.
@@
    For any prime $r>73$ such that $r\equiv 1$ (mod 4), or $r\equiv\pm 1$ (mod 5), the equation (4.8) with $c=3s$ has no non-trivial solutions with $\gcd(x,y)=1$.
@@
Let $c=2^{\alpha}\cdot 5^{\gamma}$ with $2\leq\alpha\leq 4$ and $0\leq\gamma\leq 4$ and let $r$ be a prime satisfying $r\geq 7$. Then, the equation (4.8) has no non-trivial coprime integer solutions.
Billerey and Dieulefait [23] studied the cases of $c=7$ and $c=13$ in (4.8) and proved that the following results.
@@
If either (i) $c=7$ and $r\geq 13$ or (ii) $c=13$ and $r\geq 19$, then equation (4.8) does not have any non-trivial primitive integer solutions.
In the same paper [23], the authors also proved that the equation (4.8) with $c=2^{\alpha}3^{\beta}5^{\gamma}$, that is,
@@
None of the theorems above include the case when $c$ is divisible by $11$ in (4.8). Noubissie [107] considered the equation (4.8) with $c=2^{\alpha}11$, that is,
@@
Let $r>19$ be prime and $\alpha\geq 2$. Then the equation (4.12) has no non-trivial primitive integer solutions.
@@
Let us now discuss equation (4.2) with $p>5$. In case $p=6$, Kraus [91] proved the following result.
@@
reduces to (4.2) by the trivial change of variables $y\to-y$. For even $p$, these equations are not equivalent, and (4.13) must be considered separately. Kraus [91] studied equation (4.2) with $p$ being an integer in the range $6\leq p\leq 12$. To state the results, we must first introduce the following notation.
Given an integer $N\geq 1$, we denote by ${\cal{D}}_{N}$ the set of integers $A\geq 1$ which satisfy both conditions
@@
    Then, for any integer $c$ belonging to ${\cal{D}}_{p}$, equation (4.13) has no non-trivial primitive integer solutions.
@@
    Suppose we have $p=5$ or $p=7$. Let $c$ be an integer divisible by $p$ and belonging to ${\cal{D}}_{p}$. Then, the equation (4.13) with $r=2$ has no non-trivial primitive integer solutions.
@@
We now return to equation (4.2), and next discuss it with $p=7$. Let $s$ be an integer only divisible by primes $s_{i}\not\equiv 0,1$ modulo $7$. In 2015, Freitas [68] proved that if either (i) $\alpha\geq 2$, $\beta\geq 0$ and $\gamma\geq 0$, or (ii) $\alpha=1$, $\beta\geq 1$ and $\gamma\geq 0$, or (iii) $\alpha=0$, $\beta\geq 0$ and $\gamma\geq 1$, then the equation (4.2) with $p=7$, $r\geq 17$ and $c=2^{\alpha}3^{\beta}5^{\gamma}s$ has no integer solutions where $z$ is divisible by $7$ and $|xyz|>1$. In the same paper, Freitas [68] also proved the following result.
@@
If $\alpha\geq 2$ or $\beta>0$ or $\gamma>0$, then the equation (4.2) with $p=14$, $r\geq 17$ and $c=2^{\alpha}3^{\beta}5^{\gamma}s$, (where $s$ is an integer only divisible by primes $s_{i}\not\equiv 0,1$ modulo $7$), has no primitive integer solutions with $|xyz|>1$.
In 2024, the authors of [22] use the Darmon program [51] to solve equation (4.2) for $(p,c)=(7,3)$.
@@
has no non-trivial primitive solutions.
@@
    The equation (4.2) with $p=13$ and $c=s\gamma$ has no primitive solutions $(x,y,z)$ such that $13\nmid z$ and $|xyz|>1$.
@@
    The equation (4.2) with $p=26$ and $c=10\gamma$ has no primitive solutions with $|xyz|>1$.
Billerey, Chen, Dieulefait, and Freitas [21] proved that the equation (4.2) with $c=3$, $p=13$ and $r\geq 2$ where $r$ is not a power of $7$ has no non-trivial primitive integer solutions. In a later paper, Billerey, Chen, Dembélé, Dieulefait, and Freitas [19], extended this result to all integers $r\geq 2$.
@@
4.1.2 Equation (4.2) with fixed $r$
We now consider equation (4.2) with $r$ fixed. We start with the case $r=2$, that is, equation
@@
In 1977, Terjanian [140] proved for an odd prime $p$ and integer $c\geq 2$ with no square factor or a prime divisor of the form $12kp+$ 1, if $(x,y,z)$ is a non-trivial primitive solution to (4.16) with even $z$, then $z$ must be divisible by $p$.
In 2011, Söderlund and Aldén [135] studied equation (4.16) with $p=5$ and proved the following result.
@@
Theorem 4.26 does not include the case $\alpha=1$. In the follow-up work [136], the same authors included this case as well, and thus solved equation (4.17) for $c=10$ and $c=2s$ where $s$ is a prime equal to $3$ or $7$ modulo $10$.
Bennett and Skinner [13] considered the equation (4.16) with $p\geq 4$, and they obtained the following result.
@@
If $p\geq 4$ and $c\in\{2,3,5,6,10,11,13,17\}$ then the equation (4.16) has no solutions in non-zero pairwise coprime integers $(x,y,z)$ with $x>y$, unless $(p,c)=(4,17)$ or
@@
Equation (4.16) has also been studied by Bennett and Mulholland [12], who proved that if $p$ and $c\geq 5$ are primes with $c\neq 7$ and $p>c^{12c^{2}}$, then (4.16) has no primitive solutions with $|xy|>1$ and $z$ even. They also considered the case $c=2s$ where $s\geq 5$ is prime, and proved the following theorem.
@@
has no primitive solutions with $|xy|>1$.
Bennett and Mulholland [104] studied equation (4.16) with coefficient $c=2^{\alpha}s$ where $s$ is prime and $\alpha\geq 1$, and proved the following theorem.
@@
Let $s$ be an odd square-free positive integer with $\gcd(s,21)=1$ and $\alpha\geq 1$. If prime $p$ satisfies $p>s^{132s^{2}}$, then the equation (4.18) has no primitive integer solutions.
Ivorra [83] studied equation (4.16) for $p\in\{11,13,17\}$ and obtained the following theorem.
@@
Let $p\in\{11,13,17\}$ and let $c\geq 3$ be an integer without square factors and for any prime divisor $l$ of $c$ we have $l\not\equiv 1$ modulo $p$. Then the equation (4.16) has no non-trivial primitive integer solutions.
Bennett, Vatsal and Yazdani [14] studied equation (4.2) with $r=3$, and proved the following theorems.
@@
Krawciów [92] studied equation (4.19) where $c$ is divisible by $3$ and obtained the following result.
@@
Let $p$ be a prime number, and let $c$ be a non-zero cube-free integer which is divisible by $3$. Let $C=\mathrm{rad}(c)$, then if $p>C^{10C^{2}}$, the equation (4.19) has no non-trivial primitive integer solutions.

### v2

@@ 4.1 The case $a=b=1$
We start with the case $a=b=1$. In this instance Eq. (4.1) reduces to
@@
We remark that for Eq. (4.2) Proposition 1.4 is applicable if and only if $c$ is not divisible by a $p$th power of any prime. If this condition is satisfied, we may talk about solutions in pairwise coprime integers $(x,y,z)$. From symmetry, we may also assume that $x>y$.
In 2023, Freitas and Najman [73] considered Eq. (4.2) where $p,r>3$ are primes and $c$ is an odd positive integer, not an $r$th power and $r\nmid c$. They obtained that for a set of prime exponents $p$ of positive density, the equation has no non-trivial primitive solutions $(x,y,z)$ such that $2\mid x+y$ or $r\mid x+y$.
In 2024, Kara, Mocanu and Özman [87] prove some results conditional on the weak Frey–Mazur conjecture [87, Conjecture 3.9] and the Eichler–Shimura conjecture [76, 87, Conjecture 3.11].
@@
    If $p\equiv 3$ modulo $4$, then there exists a (non-effective) constant ${\mathcal{B}}_{p}$ such that for any prime $r$ with $r>{\mathcal{B}}_{p}$, Eq. (4.2) has no non-trivial primitive integer solutions.
@@
    If $p\equiv 1$ modulo $4$, then there exists a (non-effective) constant ${\mathcal{B}}^{\prime}_{p}$ such that for any prime $r$ with $r>{\mathcal{B}}^{\prime}_{p}$, Eq. (4.2) has no non-trivial primitive integer solutions $(x,y,z)$ where $p|z$.
@@
Let $p>5$ be a fixed prime such that $p\equiv 1$ modulo $4$ and $c$ be a non-zero integer with $\gcd(c,p)=1$. Assume the Weak Frey-Mazur Conjecture and the Eichler-Shimura Conjecture. Then there exists a (non-effective) constant ${\mathcal{B}}^{\prime\prime}_{p}$ such that for any prime $r$ with $r>{\mathcal{B}}^{\prime\prime}_{p}$, Eq. (4.2) has no non-trivial primitive integer solutions.
4.1.1 Eq. (4.2) with fixed $p$
In this section, we consider Eq. (4.2) with $p\geq 3$ fixed and $r$ varies. In this section, we will not assume that $p$ is a prime. We start with the case $p=3$.
Eq. (4.2) with $p=3$, $c=2$ and $r$ even was solved by Zhang and Bai [153].
@@
Eq. (4.2) with $p=3$, $c=s^{\alpha}$ with $s$ prime, $\alpha\geq 1$ and $r\geq 4$, that is,
@@
There is an explicit set $T$ of primes containing a positive proportion of all primes, [^18] Specifically, $T$ is the set of primes $s$ for which there does not exist an elliptic curve with rational 2-torsion and conductor $2^{M}3^{2}s$, $1\leq M\leq 3$. such that for every $s\in T$, every integer $\alpha\geq 1$, and every prime $r$ satisfying $r\geq s^{8s}$ and $r\nmid\alpha$, Eq. (4.3) has no solutions in coprime non-zero integers $x,y,z$.
@@
There exists a set ${\mathcal{W}}_{3}$ of primes, with #$\#\{s\leq x:s\in{\mathcal{W}}_{3}\}\leq K\sqrt{x}\log^{2}(x)$ for some $K>0$, such that for every prime $s\geq 5$ with $s\not\in{\mathcal{W}}_{3}$, any $\alpha\geq 1$, and every prime $r$ satisfying $r\geq s^{2s}$, Eq. (4.3) has no solutions in coprime non-zero integers $x,y,z$.
Because the number of all primes up to $x$ is about $\frac{x}{\log x}$, set ${\mathcal{W}}_{3}$ contains a negligible part of all primes, hence Theorem 4.5 solves Eq. (4.3) for almost all primes $s$. There is an algorithm for checking whether any particular prime belongs to ${\mathcal{W}}_{3}$. For example, it is known [12] that all primes $s$ equal to $317$ or $1757$ modulo $2040$ do not belong to ${\mathcal{W}}_{3}$, hence Theorem 4.5 is applicable to all such primes.
Theorem 4.5 is not applicable for primes $s\in{\mathcal{W}}_{3}$. In 2018, Bennett, Bruni and Freitas [6] proved a more general version of this theorem with a smaller exceptional set.
@@
For every prime $s\geq 5$ such that $s\not\in S_{0}$, any $\alpha\geq 1$, and any prime $r$ satisfying $r\geq s^{2s}$, Eq. (4.3) has no solutions in coprime non-zero integers $x,y,z$.
As mentioned in [6], $S_{0}\subset{\mathcal{W}}_{3}$, hence Theorem 4.6 is a generalization of Theorem 4.5. The smallest examples of primes in ${\mathcal{W}}_{3}\setminus S_{0}$ are
$s=53,83,149,167,173,199,223,227,233,263,281,293,311,347,353,359,\dots$
@@
where $s$ and $r$ are primes. For small primes $s\in S_{0}$, they determined [6, Theorem 7.2] the modularity conditions on $r$ such that for $r\geq s^{2s}$ the equation has no primitive integer solutions, see Table 4.1.
Table 4.1: Modularity conditions on $r$ for small primes $s$, such that Eq. (4.5) has no primitive integer solutions if $r\geq s^{2s}$.
@@
5 13,19 or 23 (mod 24) 47 $5,11,13,17,19$ or 23 (mod 24)
11 $13,17,19$ or 23 (mod 24) 59 $5,7,11,13,19$ or 23 (mod 24)
13 11 (mod 12) 67 $7,11,13,29,37,41,43,59,$
$67,71,89,101$ or 103 (mod 120)
17 5,17 or 23 (mod 24) 71 5 (mod 6)
23 19 or 23 (mod 24) 73 41,71 or 89 (mod 120)
29 $7,11,13,17,19$ or 23 (mod 24) 79 $5,7,11,13,19$ or 23 (mod 24)
31 5 or 11 (mod 24) 89 $13,17,19$ or 23 (mod 24)
41 $5,7,11,17,19$ or 23 (mod 24) 97 11 (mod 12)
@@
If $r$ is prime with $r\equiv 13,19$ or $23$ (mod $24$), then there are no non-trivial primitive integer solutions $(x,y,z)$ to the equation
@@
Set $S_{0}$ contains all primes of the form $s=2^{a}3^{b}-1$ for integers $a\geq 5$ and $b\geq 1$, hence these primes are not covered by Theorem 4.6. For such primes, we have the following result [6], which uses a Jacobi symbol [^19] For the definition and its main properties, see [45, Section 1.4.2]. condition.
@@
then Eq. (4.3) has no non-trivial primitive solutions.
@@
We next discuss Eq. (4.2) with $p=4$ and $c,r$ primes, that is
@@
We first discuss Eq. (4.7) with $c=73,89$ or $113$ and $r>13$. Dieulefait [61] proved the following theorem.
@@
If $r$ is a prime with $r>13$, then Eq. (4.7) with $c=73,89$ or $113$ does not have any non-trivial primitive integer solutions.
@@
We next discuss Eq. (4.2) with $p=5$, that is,
@@
There exists a set ${\mathcal{W}}_{5}$ of primes, with #$\#\{s\leq x:s\in{\mathcal{W}}_{5}\}\leq K\sqrt{x}\log^{2}(x)$ for some $K>0$, such that for every prime $s\geq 5$ such that $s\not\in{\mathcal{W}}_{5}$, any $\alpha\geq 1$, and any prime $r$ satisfying $r\geq s^{13s}$, equation
@@
Bruni [34] also developed methods for solving Eq. (4.9) for certain primes $s$ belonging to ${\mathcal{W}}_{5}$. In fact, he solved Eq. (4.9) for all primes $s$ outside the specific families listed in Table 7.6 of [34].
For Eq. (4.8) with $c=2$, that is,
@@
For all primes $r$, Eq. (4.10) has no non-trivial primitive integer solutions $(x,y,z)$ such that $z$ is divisible by either $2$, $5$, or $r$.
Eq. (4.8) with $c=3$, that is,
@@
Eq. (4.11) has no non-trivial primitive integer solutions for any integer $r\geq 2$.
We next discuss Eq. (4.8) for other values of $c$. In 2013, Dieulefait and Freitas [63] proved the following theorem.
@@
    For any prime $r>13$ such that $r\equiv 1$ (mod 4), or $r\equiv\pm 1$ (mod 5), Eq. (4.8) with $c=2s$ has no non-trivial integer solutions with $\gcd(x,y)=1$.
@@
    For any prime $r>73$ such that $r\equiv 1$ (mod 4), or $r\equiv\pm 1$ (mod 5), Eq. (4.8) with $c=3s$ has no non-trivial integer solutions with $\gcd(x,y)=1$.
@@
Let $c=2^{\alpha}\cdot 5^{\gamma}$ with $2\leq\alpha\leq 4$ and $0\leq\gamma\leq 4$ and let $r$ be a prime satisfying $r\geq 7$. Then, Eq. (4.8) has no non-trivial coprime integer solutions.
Billerey and Dieulefait [24] studied the cases of $c=7$ and $c=13$ in (4.8) and proved the following results.
@@
If either (i) $c=7$ and $r\geq 13$ or (ii) $c=13$ and $r\geq 19$, then Eq. (4.8) does not have any non-trivial primitive integer solutions.
In the same paper [24], the authors also proved that Eq. (4.8) with $c=2^{\alpha}3^{\beta}5^{\gamma}$, that is,
@@
None of the theorems above include the case when $c$ is divisible by $11$ in (4.8). Noubissie [110] considered Eq. (4.8) with $c=2^{\alpha}11$, that is,
@@
Let $r>19$ be prime and $\alpha\geq 2$. Then Eq. (4.12) has no non-trivial primitive integer solutions.
@@
Let us now discuss Eq. (4.2) with $p>5$. In case $p=6$, Kraus [93] proved the following result.
@@
reduces to (4.2) by the trivial change of variables $y\to-y$. For even $p$, these equations are not equivalent, and (4.13) must be considered separately. Kraus [93] studied Eq. (4.13) with $p$ being an integer in the range $6\leq p\leq 12$. To state the results, we must first introduce the following notation.
Given an integer $N\geq 1$, we denote by ${\mathcal{D}}_{N}$ the set of integers $A\geq 1$ which satisfy both conditions
@@
    Then, for any integer $c$ belonging to ${\mathcal{D}}_{p}$, Eq. (4.13) has no non-trivial primitive integer solutions.
@@
    Suppose we have $p=5$ or $p=7$. Let $c$ be an integer divisible by $p$ and belonging to ${\mathcal{D}}_{p}$. Then, Eq. (4.13) with $r=2$ has no non-trivial primitive integer solutions.
@@
We now return to Eq. (4.2), and next discuss it with $p=7$. Let $s$ be an integer only divisible by primes $s_{i}\not\equiv 0,1$ modulo $7$. In 2015, Freitas [70] proved that if either (i) $\alpha\geq 2$, $\beta\geq 0$ and $\gamma\geq 0$, or (ii) $\alpha=1$, $\beta\geq 1$ and $\gamma\geq 0$, or (iii) $\alpha=0$, $\beta\geq 0$ and $\gamma\geq 1$, then Eq. (4.2) with $p=7$, $r\geq 17$ and $c=2^{\alpha}3^{\beta}5^{\gamma}s$ has no integer solutions where $z$ is divisible by $7$ and $|xyz|>1$. In the same paper, Freitas [70] also proved the following result.
@@
If $\alpha\geq 2$ or $\beta>0$ or $\gamma>0$, then Eq. (4.2) with $p=14$, $r\geq 17$ and $c=2^{\alpha}3^{\beta}5^{\gamma}s$, (where $s$ is an integer only divisible by primes $s_{i}\not\equiv 0,1$ modulo $7$), has no primitive integer solutions with $|xyz|>1$.
In 2024, the authors of [23] use the Darmon program [53] to solve Eq. (4.2) for $(p,c)=(7,3)$.
@@
has no non-trivial primitive integer solutions.
@@
    Eq. (4.2) with $p=13$ and $c=s\gamma$ has no primitive integer solutions $(x,y,z)$ such that $13\nmid z$ and $|xyz|>1$.
@@
    Eq. (4.2) with $p=26$ and $c=10\gamma$ has no primitive integer solutions with $|xyz|>1$.
Billerey, Chen, Dieulefait, and Freitas [22] proved that Eq. (4.2) with $c=3$, $p=13$ and $r\geq 2$ where $r$ is not a power of $7$ has no non-trivial primitive integer solutions. In a later paper, Billerey, Chen, Dembélé, Dieulefait, and Freitas [20] extended this result to all integers $r\geq 2$.
@@
4.1.2 Eq. (4.2) with fixed $r$
We now consider Eq. (4.2) with $r$ fixed. We start with the case $r=2$, that is, equation
@@
In 1977, Terjanian [145] proved for an odd prime $p$ and integer $c\geq 2$ with no square factor or a prime divisor of the form $12kp+$ 1, if $(x,y,z)$ is a non-trivial primitive integer solution to (4.15) with even $z$, then $z$ must be divisible by $p$.
In 2011, Söderlund and Aldén [140] studied Eq. (4.15) with $p=5$ and proved the following result.
@@
Theorem 4.26 does not include the case $\alpha=1$. In the follow-up work [141], the same authors included this case as well, and thus solved Eq. (4.16) for $c=10$ and $c=2s$ where $s$ is a prime equal to $3$ or $7$ modulo $10$.
Bennett and Skinner [15] considered Eq. (4.15) with $p\geq 4$, and they obtained the following result.
@@
If $p\geq 4$ and $c\in\{2,3,5,6,10,11,13,17\}$ then Eq. (4.15) has no solutions in non-zero pairwise coprime integers $(x,y,z)$ with $x>y$, unless $(p,c)=(4,17)$ or
@@
Eq. (4.15) has also been studied by Bennett and Mulholland [14], who proved that if $p$ and $c\geq 5$ are primes with $c\neq 7$ and $p>c^{12c^{2}}$, then (4.15) has no primitive solutions with $|xy|>1$ and $z$ even. They also considered the case $c=2s$ where $s\geq 5$ is prime, and proved the following theorem.
@@
has no primitive integer solutions with $|xy|>1$.
Bennett and Mulholland [107] studied Eq. (4.15) with coefficient $c=2^{\alpha}s$ where $s$ is prime and $\alpha\geq 1$, and proved the following theorem.
@@
Let $s$ be an odd square-free positive integer with $\gcd(s,21)=1$ and $\alpha\geq 1$. If prime $p$ satisfies $p>s^{132s^{2}}$, then Eq. (4.17) has no primitive integer solutions.
Ivorra [85] studied Eq. (4.15) for $p\in\{11,13,17\}$ and obtained the following theorem.
@@
Let $p\in\{11,13,17\}$ and let $c\geq 3$ be an integer without square factors and for any prime divisor $l$ of $c$ we have $l\not\equiv 1$ modulo $p$. Then Eq. (4.15) has no non-trivial primitive integer solutions.
Bennett, Vatsal and Yazdani [16] studied Eq. (4.2) with $r=3$, and proved the following theorems.
@@
Krawciów [94] studied Eq. (4.18) where $c$ is divisible by $3$ and obtained the following result.
@@
Let $p$ be a prime number, and let $c$ be a non-zero cube-free integer which is divisible by $3$. Let $C=\mathrm{rad}(c)$, then if $p>C^{10C^{2}}$, Eq. (4.18) has no non-trivial primitive integer solutions.